# Graph Kernel Neural Networks

### v2

@@ III-F Computational complexity and runtime analysis
In terms of runtime, for the datasets considered in this paper, this translates to the results shown in Table I, where we report the average epoch training time and the average inference time per dataset obtained during the grid search (for the best model, for each dataset and fold). Both training and testing of each model were performed using an Intel(R) Xeon(R) Silver 4114 CPU @ 2.20GHz processor and an NVIDIA Tesla V100. Note that in order to exploit the computational power of GPUs better we implemented a GPU version of the WL kernel. Although the resulting inference runtime is higher than three widely used baselines, we compared our model to (ECC [53], DiffPool [54], and GIN [32]), it is of a similar order of magnitude. When it comes to training, on the other hand, the computational bottleneck of our network is the computation of the numerical gradient, which in turn explains the subpar runtime.
@@ III-G Relation to existing works
The closest works to ours we are aware of are [13, 14]. Nikolentzos et al. propose a GNN where the first layer consists of a series of hidden graphs that are compared against the input graph using a random walk kernel [13]. However, due to their optimization strategy, their model only works with a single differentiable kernel, whereas our model allows us to tap into the expressive power of any type of kernel. Moreover, in the architecture of [13], the learned structural masks are assumed to be complete weighted graphs, whereas we allow for graphs with arbitrary structure. Finally, note that [14] works very much in the same way of [13], as it relies on the underlying graph kernel to be differentiable and, more specifically, it makes use of the random walk graph kernel. Another similarity with [13] is that the model proposed in [14] is also not purely structural, as opposed to ours. Instead, it learns matrices of weights which, in order to be interpreted as learned structural patterns, need to be first thresholded using a ReLU function that prunes edges associated with smaller weights.
@@ IV-A Datasets
MUTAG [55] is a dataset consisting of 188 mutagenic aromatic and heteroaromatic nitro compounds (with 7 discrete labels) assayed for mutagenicity in Salmonella typhimurium. The goal is predicting whether each compound possesses mutagenicity effect on the Gram-negative bacterium Salmonella typhimurium.
NCI1 (the anti-cancer activity prediction dataset) is a balanced subset of datasets of chemical compounds screened for activity against non-small cell lung cancer and ovarian cancer cell lines [56]. It has 37 discrete labels.
The PROTEINS [57, 58] dataset consists of proteins represented as graphs. Nodes represent the amino acids and an edge occurs between two vertices if they are neighbors in the amino-acid sequence or 3D space. It has 3 discrete labels, representing helix, sheet, or turn. The aim is to distinguish proteins into enzymes and non-enzymes.
The PTC (Predictive Toxicology Challenge) dataset [59, 60] is a collection of chemical compounds reporting the carcinogenicity for Male Rats (MR), Female Rats (FR), Male Mice (MM), and Female Mice (FM). We selected graphs of Male Rats (MR) for evaluation, consisting of 344 graphs very small and sparse with 19 discrete labels.
ENZYMES consists of protein tertiary structures obtained from [57], in particular 600 enzymes from the BRENDA enzyme database [61]. The task for this dataset is assigning each enzyme to one of the 6 EC top-level classes. For our experiments, we considered only 3 discrete labels.
@@
The IMDB-BINARY/MULTI datasets [62] collect graphs about movie collaboration. Each graph consists of nodes representing actors/actresses who played roles in movies in IMDB, and an edge between two vertices takes place if they appear in the same movie. Graphs are derived from the Action and Romance genres. The task is to identify which genre an ego-network graph belongs to. Note that while the bio/chemo-informatics graphs described above have categorical features, the nodes have no features in the IMDB-BINARY/MULTI datasets.
For the graph regression task, we employ a common benchmark molecular dataset, ZINC [63], a drug-constrained solubility prediction dataset. In particular, we use a subset (12K) of the ZINC molecular graphs (250K), as in [64]. Here, the goal is regressing the constrained solubility, which is the desired molecular property. The node features in each molecular graph refer to the types of heavy atoms. Even though the dataset takes into consideration the types of bonds between nodes encoded as edge features, we do not use them in our experiments.
@@ IV-C Interpretability analysis
One of the factors that fostered interpretability in classical CNNs is the possibility to visualize and analyze the learned filters, capturing the fundamental structures characterizing the input images [13]. This feature is missing in the classical formulation of graph convolution relying on the message-passing paradigm. On the other hand, our method learns actual graph masks, potentially increasing the interpretability of the model by allowing us to probe into the explanatory factors of variation behind the input data through the learned structural masks.
@@ IV-E Graph classification and regression
We compare our model against: 6 state-of-the-art GNNs: ECC [53], DGCNN [46], DiffPool [54], GIN [32], (s)GIN [65], and GraphSAGE [66]; two distinct baselines, depending on the dataset type (see below): Molecular Fingerprint [67, 68] and Deep Multisets [69]; the WL kernel [7]; RWNN [13] and KerGNN [14], two GNN models employing differentiable graph kernels, the closest existing neural architectures to ours. We use a $C$-SVM [70] classifier for the WL subtree kernel. For the baselines, we follow [71] and use the Molecular Fingerprint technique [67, 68] for chemical datasets. For the social dataset, we rely on the permutation invariant model of [69].
@@
We perform a grid search for all the methods to optimize the hyper-parameters. In particular, for the WL method, we optimize the value of $C$ and the number of WL iterations $h\in\{4,5,6,7\}$. For the RWGNN, we investigate the hyper-parameter ranges used by the authors [13], while for all the other GNNs and the baselines, we follow [71] and we perform a full search over the hyper-parameters grid. For our model, we explore the following hyper-parameters: number of structural masks in $\{8,16,32\}$, maximum number of nodes of a structural mask in $\{6,8,10\}$, subgraph radius in $\{1,2,3\}$, and number of layers in $\{1,2,3\}$. We implemented two versions of our network, one that makes use of the WL kernel [7] and one that makes use of the Graphlet (GL) kernel [17] (the top 2 performing kernels identified in the ablation study). For the GL kernel, we use graphlets of size 5 on all datasets except for IMDB-B and IMDB-M, where we limit the size to 3222This is due to the presence of high degree nodes in the IMDB datasets and the fact that the computational complexity of the GL kernel is $O(Nd_{\max}^{k-1})$, where $N$ is the number of nodes of the graph and $d_{\max}$ is the maximum degree..
@@
Results. The accuracy for each method and dataset is reported in Table II. The performance of our approach outperforms the baselines [67, 69, 68] on all the datasets, with the improvement being statistically significant on all datasets except PTC. Compared to other GNNs, the performance of our model is on par with the best ones, as we can see for MUTAG, PROTEINS, and PTC. This holds as long as the number of node labels is small. Indeed, in the case of high-dimensional node labels (e.g., NCI1), the classification accuracy tends to decrease as the optimization becomes harder. As for the social dataset, the performance of our approach is satisfactory, particularly considering the lack of discriminative sub-structures in social networks. Finally, our method achieves the best performance (i.e., lowest MAE) on ZINC, with only GIN achieving a similar yet higher MAE. It should also be noted that our model always outperforms or performs comparably to both KerGNN and RWGNN, the only two other models that allow some degree of interpretability of the learned structures.
@@ Acknowledgments
L.C. and A.B. are supported by the PRIN 2022 project n. 2022AL45R2 (EYE-FI.AI, CUP H53D2300350-0001). G.M. acknowledges financial support from the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS__\__00000043 – CUP H43C22000540006). In this regard, the views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.

### v3

@@ III-F Computational complexity and runtime analysis
In terms of runtime, for the datasets considered in this paper, this translates to the results shown in Table I, where we report the average epoch training time and the average inference time per dataset obtained during the grid search (for the best model, for each dataset and fold). Both training and testing of each model were performed using an Intel(R) Xeon(R) Silver 4114 CPU @ 2.20GHz processor and an NVIDIA Tesla V100. Note that in order to exploit the computational power of GPUs better we implemented a GPU version of the WL kernel. Although the resulting inference runtime is higher than three widely used baselines, we compared our model to (ECC [simonovsky2017dynamic], DiffPool [ying2018hierarchical], and GIN [xu2018powerful]), it is of a similar order of magnitude. When it comes to training, on the other hand, the computational bottleneck of our network is the computation of the numerical gradient, which in turn explains the subpar runtime.
@@ III-G Relation to existing works
The closest works to ours we are aware of are [nikolentzos2020random, feng2022kergnns]. Nikolentzos et al. propose a GNN where the first layer consists of a series of hidden graphs that are compared against the input graph using a random walk kernel [nikolentzos2020random]. However, due to their optimization strategy, their model only works with a single differentiable kernel, whereas our model allows us to tap into the expressive power of any type of kernel. Moreover, in the architecture of [nikolentzos2020random], the learned structural masks are assumed to be complete weighted graphs, whereas we allow for graphs with arbitrary structure. Finally, note that [feng2022kergnns] works very much in the same way of [nikolentzos2020random], as it relies on the underlying graph kernel to be differentiable and, more specifically, it makes use of the random walk graph kernel. Another similarity with [nikolentzos2020random] is that the model proposed in [feng2022kergnns] is also not purely structural, as opposed to ours. Instead, it learns matrices of weights which, in order to be interpreted as learned structural patterns, need to be first thresholded using a ReLU function that prunes edges associated with smaller weights.
@@ IV-A Datasets
MUTAG [debnath1991structure] is a dataset consisting of 188 mutagenic aromatic and heteroaromatic nitro compounds (with 7 discrete labels) assayed for mutagenicity in Salmonella typhimurium. The goal is predicting whether each compound possesses mutagenicity effect on the Gram-negative bacterium Salmonella typhimurium.
NCI1 (the anti-cancer activity prediction dataset) is a balanced subset of datasets of chemical compounds screened for activity against non-small cell lung cancer and ovarian cancer cell lines [wale2008comparison]. It has 37 discrete labels.
The PROTEINS [borgwardt2005protein, dobson2003distinguishing] dataset consists of proteins represented as graphs. Nodes represent the amino acids and an edge occurs between two vertices if they are neighbors in the amino-acid sequence or 3D space. It has 3 discrete labels, representing helix, sheet, or turn. The aim is to distinguish proteins into enzymes and non-enzymes.
The PTC (Predictive Toxicology Challenge) dataset [helma2001predictive, kriege2012subgraph] is a collection of chemical compounds reporting the carcinogenicity for Male Rats (MR), Female Rats (FR), Male Mice (MM), and Female Mice (FM). We selected graphs of Male Rats (MR) for evaluation, consisting of 344 graphs very small and sparse with 19 discrete labels.
ENZYMES consists of protein tertiary structures obtained from [borgwardt2005protein], in particular 600 enzymes from the BRENDA enzyme database [schomburg2004brenda]. The task for this dataset is assigning each enzyme to one of the 6 EC top-level classes. For our experiments, we considered only 3 discrete labels.
@@
The IMDB-BINARY/MULTI datasets [yanardag2015deep] collect graphs about movie collaboration. Each graph consists of nodes representing actors/actresses who played roles in movies in IMDB, and an edge between two vertices takes place if they appear in the same movie. Graphs are derived from the Action and Romance genres. The task is to identify which genre an ego-network graph belongs to. Note that while the bio/chemo-informatics graphs described above have categorical features, the nodes have no features in the IMDB-BINARY/MULTI datasets.
For the graph regression task, we employ a common benchmark molecular dataset, ZINC [irwin2012zinc], a drug-constrained solubility prediction dataset. In particular, we use a subset (12K) of the ZINC molecular graphs (250K), as in [dwivedi2020benchmarking]. Here, the goal is regressing the constrained solubility, which is the desired molecular property. The node features in each molecular graph refer to the types of heavy atoms. Even though the dataset takes into consideration the types of bonds between nodes encoded as edge features, we do not use them in our experiments.
@@ IV-C Interpretability analysis
One of the factors that fostered interpretability in classical CNNs is the possibility to visualize and analyze the learned filters, capturing the fundamental structures characterizing the input images [nikolentzos2020random]. This feature is missing in the classical formulation of graph convolution relying on the message-passing paradigm. On the other hand, our method learns actual graph masks, potentially increasing the interpretability of the model by allowing us to probe into the explanatory factors of variation behind the input data through the learned structural masks.
@@ IV-E Graph classification and regression
We compare our model against: 6 state-of-the-art GNNs: ECC [simonovsky2017dynamic], DGCNN [zhang2018end], DiffPool [ying2018hierarchical], GIN [xu2018powerful], (s)GIN [di2020mutual], and GraphSAGE [hamilton2017inductive]; two distinct baselines, depending on the dataset type (see below): Molecular Fingerprint [ralaivola2005graph, luzhnica2019graph] and Deep Multisets [zaheer2017deep]; the WL kernel [shervashidze2011weisfeiler]; RWNN [nikolentzos2020random] and KerGNN [feng2022kergnns], two GNN models employing differentiable graph kernels, the closest existing neural architectures to ours. We use a $C$-SVM [chang2011libsvm] classifier for the WL subtree kernel. For the baselines, we follow [errica_fair_2020] and use the Molecular Fingerprint technique [ralaivola2005graph, luzhnica2019graph] for chemical datasets. For the social dataset, we rely on the permutation invariant model of [zaheer2017deep].
@@
We perform a grid search for all the methods to optimize the hyper-parameters. In particular, for the WL method, we optimize the value of $C$ and the number of WL iterations $h\in\{4,5,6,7\}$. For the RWGNN, we investigate the hyper-parameter ranges used by the authors [nikolentzos2020random], while for all the other GNNs and the baselines, we follow [errica_fair_2020] and we perform a full search over the hyper-parameters grid. For our model, we explore the following hyper-parameters: number of structural masks in $\{8,16,32\}$, maximum number of nodes of a structural mask in $\{6,8,10\}$, subgraph radius in $\{1,2,3\}$, and number of layers in $\{1,2,3\}$. We implemented two versions of our network, one that makes use of the WL kernel [shervashidze2011weisfeiler] and one that makes use of the Graphlet (GL) kernel [shervashidze2009efficient] (the top 2 performing kernels identified in the ablation study). For the GL kernel, we use graphlets of size 5 on all datasets except for IMDB-B and IMDB-M, where we limit the size to 3222This is due to the presence of high degree nodes in the IMDB datasets and the fact that the computational complexity of the GL kernel is $O(Nd_{\max}^{k-1})$, where $N$ is the number of nodes of the graph and $d_{\max}$ is the maximum degree..
@@
Results. The accuracy for each method and dataset is reported in Table II. The performance of our approach outperforms the baselines [ralaivola2005graph, zaheer2017deep, luzhnica2019graph] on all the datasets, with the improvement being statistically significant on all datasets except PTC. Compared to other GNNs, the performance of our model is on par with the best ones, as we can see for MUTAG, PROTEINS, and PTC. This holds as long as the number of node labels is small. Indeed, in the case of high-dimensional node labels (e.g., NCI1), the classification accuracy tends to decrease as the optimization becomes harder. As for the social dataset, the performance of our approach is satisfactory, particularly considering the lack of discriminative sub-structures in social networks. Finally, our method achieves the best performance (i.e., lowest MAE) on ZINC, with only GIN achieving a similar yet higher MAE. It should also be noted that our model always outperforms or performs comparably to both KerGNN and RWGNN, the only two other models that allow some degree of interpretability of the learned structures.
@@ Acknowledgments
L.C. and A.B. are supported by the PRIN 2022 project n. 2022AL45R2 (EYE-FI.AI, CUP H53D2300350-0001). G.M. acknowledges financial support from the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS_\_00000043 – CUP H43C22000540006). In this regard, the views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.
[Uncaptioned image] Luca Cosmo is an Assistant Professor in Computer Science at University Ca’ Foscari of Venice, working in Geometric Deep Learning and Generative AI, with more than 50 published papers. He has been serving on the program committee and as a reviewer for top conferences and journals in these fields and he gave short courses and tutorials in geometry processing related topics at major conferences in the field. He is a member of ELLIS and co-founded a start-up working on geometry processing for industrial product inspection.
[Uncaptioned image] Giorgia Minello Giorgia Minello worked in the industry for almost ten years before receiving, in 2019, her PhD in Computer Science from Ca’ Foscari, University of Venice (Italy), where she currently holds a Postdoctoral Research Fellow position. She is also a scientific collaborator for the IESE Business School, University of Navarra, in the context of a social network analysis project. Her research mainly focuses on Structural Pattern Recognition, Machine Learning and Natural Language Processing. The leading topic of her studies concerns the analysis of networks, such as the use of quantum approaches for structural analysis or the application of network analysis methods for mining textual corpora.
[Uncaptioned image] Alessandro Bicciato received the PhD in computer science from the Ca’ Foscari University of Venice, Italy in 2024. He is currently working as a post-doctorate researcher at the Computer Science Department of Ca’ Foscari University of Venice, Italy. His research interests include network science and pattern recognition.
Michael Bronstein Michael Bronstein is the DeepMind Professor of AI at the University of Oxford. He previously served as Head of Graph Learning Research at Twitter, professor at Imperial College London, and held visiting appointments at Stanford, MIT, and Harvard. He is the recipient of the Royal Society Wolfson Research Merit Award, Royal Academy of Engineering Silver Medal, Turing World-Leading AI Research Fellowship, five ERC grants, two Google Faculty Research Awards, and two Amazon AWS ML Research Awards. He is a Member of the Academia Europaea, Fellow of IEEE, IAPR, BCS, and ELLIS, ACM Distinguished Speaker, and World Economic Forum Young Scientist. In addition to his academic career, Michael is a serial entrepreneur and founder of multiple startup companies, including Novafora, Invision (acquired by Intel in 2012), Videocites, and Fabula AI (acquired by Twitter in 2019). He is the Chief Scientist at VantAI and scientific advisor at Recursion Pharmaceuticals.
Emanuele Rodolà is Professor of Computer Science at Sapienza University of Rome, where he leads the GLADIA group of Geometry, Learning and Applied AI, funded by an ERC Grant and a Google Research Award. He is currently an ELLIS fellow, and formerly fellow of the Humboldt Foundation and the Japan Society for the Promotion of Science. His research interests lie at the intersection of geometry processing, graph theory, geometric deep learning, applied AI and computer vision, and has published more than 100 papers in these areas.
[Uncaptioned image] Luca Rossi received the PhD degree in computer science from the Ca’ Foscari University of Venice, Italy, in 2013. He is currently an Assistant Professor with the Hong Kong Polytechnic University, having held various positions with the University of Birmingham, Aston University, Southern University of Science and Technology, and Queen Mary University of London. He has published more than 50 papers in international journals and conferences. His research interests include the areas of pattern recognition, data mining, and network science. He is currently a member of the editorial board of the journal Pattern Recognition and vice-chair of the technical committee 2 of the International Association for Pattern Recognition.
[Uncaptioned image] Andrea Torsello received his PhD in computer science at the University of York, UK. From 2007 he is with Ca’Foscari University of Venice, Italy, where he is Full Professor. He has published over 150 papers in international journals and conferences. His research interests are in the areas of Computer Vision and Pattern Recognition, in particular the interplay between Stochastic and Structural approaches as well as Game-Theoretic and Physical models. He is currently a member of the editorial board of the journal Pattern Recognition and chair of the technical committee 2 of the International Association for Pattern Recognition.